40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 11:58:45 2022, Paco, Summary, General, Jenne laser kaput? 
    Reply  Wed Jul 20 14:12:07 2022, Paco, Summary, General, Jenne laser kaput! 
       Reply  Wed Jul 20 15:58:52 2022, Koji, Summary, General, Jenne laser kaput! 
          Reply  Mon Jul 25 09:05:50 2022, Paco, Summary, General, Testing 950nm laser found in trash pile PXL_20220722_234600124.jpgPXL_20220722_234551918.jpg
             Reply  Fri Aug 5 17:03:31 2022, Yehonathan, Summary, General, Testing 950nm laser found in trash pile 20220805_164434.jpgLaserActuation_TF_Measurement.drawio.pdf
                Reply  Wed Sep 28 19:15:56 2022, Koji, Update, General, Testing 950nm laser found in trash pile PXL_20220929_013850989.MP.jpgPXL_20220929_013859439.jpgPXL_20220929_013911125.jpg
                   Reply  Wed Nov 30 18:27:07 2022, Paco, Update, General, Making the Jenne laser great again PXL_20221201_004623398.jpgPXL_20221201_015752792.jpg
Message ID: 17022     Entry time: Wed Jul 20 14:12:07 2022     In reply to: 17021     Reply to this: 17023
Author: Paco 
Type: Summary 
Category: General 
Subject: Jenne laser kaput! 

[Koji, Yehonathan, Paco]

Koji pointed out that this laser was always driven with a current driver (which was not nearby), and after finding it on one of the rolling carts, we hooked up the system but found that the laser driver displayed open circuit near the usual 20mA operating point. We therefore have to conclude that this laser is no more. We will look for a reasonable replacement.

Quote:

[Paco, Yehonathan, JC]

We were trying to setup the Jenne laser to characterize the response of three 1811s that Yehonathan is using for his WOPA experiment (in QIL). We hooked up a ~ 5 VDC power supply to the bias tee and looked to see if there was any DC response in the REF PD. We used a DB9 breakout board and a DB9 cable, and saw some current being drawn. The DC current was a bit too high (500 mA), so we turned the DC voltage off, and realized the VDC power was reversed, probably along the DB9 cable which we didn't check before. As we flipped the power supply leads and turned power back on, we could no longer see any current even though the voltage was now right (or was it???). We would like to debug this laser, and continue using it if it still works (!), but there is negligible documentation either here or in the wiki, so if there are any known places to look at it would be helpful to know them.

 

ELOG V3.1.3-